OpenStudy (anonymous):

Given the region bounded by the graphs y=xsin(x), y=0, x=0, x=pi, find a) the volume of the solid generated by revolving the region about the x-axis, b) the volume of the solid generated by revolving the region about the y-axis, c) the centroid of the region (assume p=1) by using disk/washer/shell methods and integration by parts.

OpenStudy (amistre64):

which part are you having issues with?

OpenStudy (anonymous):

x in the front for both

OpenStudy (amistre64):

y = xsin(x) from 0 to pi, revolved around the x axis tells us that the radius of each piece is equal to y the area of each slice of rotation is: pi r^2\[\int_{0}^{pi}~pi~(xsin(x))^2~dx\] \[pi \int_{0}^{pi}~x^2~sin^2(x)~dx\] my first thought is a by parts setup

OpenStudy (amistre64):

of using a trig identity on the sin^2 cos(2a) = cos^2(a) - sin^2(a) = (1-sin^2(a)) - sin^2(a) = 1 -2sin^2(a) cos(2a) - 1= -2sin^2(a) \[\frac12-\frac12cos(2a)=sin^2(a)\] \[\int \frac{x^2}2-\frac{x^2}2cos(2x)~dx\] might make for an easier run

OpenStudy (amistre64):

we can table this out as: -cos(2x)/2 x^2 -sin(2x)/4 | - 2x cos(2x)/8 | multiply these rows and add them together + 2 sin(2x)/16 | \[\int -\frac{x^2}2cos(2x)~dx=-\frac{x^2}{4}sin(2x)-\frac{x}{4}cos(2x)+\frac18sin(2x)\] if i didnt make an error

OpenStudy (anonymous):

so what happens to the X^2

OpenStudy (anonymous):

oh okay i got it yup thank u

OpenStudy (anonymous):

so what will you do for the y-axis

OpenStudy (amistre64):

the y axis would require us to look at the function as x = f(y), which is the inverse, but that may require some extra thought

OpenStudy (amistre64):

i believe the shell method would be better for the rotation around the y axis

OpenStudy (amistre64):

the area of the walls of a cylinder (a shell) is 2pi r from the base circumference, times the height: Area = 2pi r h in this case, r relates to x, and h relates to x sinx\[\int_{0}^{pi}~2pi~x(xsin(x))~dx\]

OpenStudy (amistre64):

ugh, i did x and not x^2 lol

OpenStudy (amistre64):

sin(x) x^2 -cos(x) - 2x -sin(x) + 2 cos(x) at 0 we get -2 at pi we get pi^2 - 2 2pi(pi^2-4)

OpenStudy (anonymous):

so what you just did is the the yaxis

OpenStudy (anonymous):

and for the centroid

OpenStudy (amistre64):

i believe centroid refers to a point that balances the whole area on the tip of a pin ... if that makes any sense. assuming the density is uniform thruout, you will want to know the total area, which is just an integration of x sinx from 0 to pi, a rather simple by parts knowing this, you can solve for half of it: \[Area:\int_{0}^{pi}f(x)~dx=F(pi)-F(0)\] \[\frac12Area:\int_{0}^{b}f(x)~dx=F(b)-F(0)\] \[F(b)=\frac12Area-F(0)\] and solve for b

OpenStudy (amistre64):

this would give you x=b as the balance point along that dimension, still pondering the y tho

OpenStudy (anonymous):

can you plz finish it

OpenStudy (amistre64):

i would perfer to see some of your work on that last explanation i did; the integration itself is pretty simple enough to do. as for the y axis, im thinking we could need to develop a strategy of subtracting a constant from the function to determine a half area

OpenStudy (amistre64):

but i cant see a good way to determine that, we would have to parse it another way, split it into 3 sections such that \[Area=\int_{0}^{c}xsin(x)~dx+\int_{c}^{d}xsin(x)~dx+\int_{d}^{pi}xsin(x)~dx\] but i havent got a real good idea for it yet

OpenStudy (anonymous):

so why did u do it 3 times

OpenStudy (amistre64):

i was just considering something that came to mind, but i cant see a way that it makes it simpler in this instance, so i have to re-examine it

OpenStudy (amistre64):

the area is equal to pi. so half of the area is pi/2 but im running into a blank solving for: sin(b) - b cos(b) = pi/2

OpenStudy (amistre64):

i got no good ideas coming to me for the centroid :/ you have any ideas?